OpenStudy (anonymous):

Write the expression as an algebraic expression in x for x>0: 4 tan(arccos x)

OpenStudy (anonymous):

\[ \cos \left(\cos ^{-1}(x)\right)=x\\ \sin \left(\cos ^{-1}(x)\right)=\sqrt{1-x^2 }\\Hence\\ \tan \left(\cos ^{-1}(x)\right)=\frac{\sqrt {1-x^2}}{x}\\ 4\tan \left(\cos ^{-1}(x)\right)=\frac{4\sqrt {1-x^2}}{x}\\ \]

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

YW